Determine the number of moles of argon gas at STP, which occupies 56.2 liters.
1
Expert's answer
2019-05-22T07:04:00-0400

Solution.

"p \times V = n \times R \times T"

"n = \frac{p \times V}{R \times T}"

p = 1.013*10^5 Pa

V = 22.4 l/mole

T = 273 K


"n = \frac{1.013 \times 10^5 \times 22.4}{8.31 \times 273}"

n = 1000.22 mole

Answer:

n = 1000.22 mole